# GARLIC: GAussian Representation LearnIng for spaCe partitioning

### v1

@@ Abstract
We introduce GARLIC (GAussian Representation LearnIng for spaCe partitioning), a novel indexing structure based on $N$-dimensional Gaussians for efficiently learning high-dimensional vector spaces. Our approach is inspired from Gaussian splatting techniques, typically used in 3D rendering, which we adapt for high-dimensional search and classification. We optimize Gaussian parameters using information-theoretic objectives that balance coverage, assignment confidence, and structural and semantic consistency. A key contribution is to progressively refine the representation through split and clone operations, handling hundreds of dimensions, thus handling varying data densities. GARLIC offers the fast building times of traditional space partitioning methods (e.g., under $\sim 5$ min build time for SIFT1M) while achieving $\sim 50\%$ Recall10@10 in low-candidate regimes. Experimental results on standard benchmarks demonstrate our method’s consistency in (a) $k$-NN retrieval, outperforming methods, such as Faiss-IVF, in fast-recall by using about half their probes for the same Recall10@10 in Fashion-MNIST, and (b) in classification tasks, beating by $\sim 15\%$ accuracy other majority voting methods. Further, we show strong generalization capabilities, maintaining high accuracy even with downsampled training data: using just $1\%$ of the training data returns $\sim 45\%$ Recall@1, thus making GARLIC quite powerful for applications requiring both speed and accuracy.
@@ 1 Introduction
Let $X=\{x_{i}\}_{i=1}^{n}\subset\mathbb{R}^{d}$ be a set of high-dimensional embeddings, derived from handcrafted descriptors (e.g., SIFT1M [37]) or learned via neural networks, which encode semantic and geometric similarity. Such embeddings are commonly used in tasks such as $k$-nearest neighbor ($k$-NN) retrieval and classification. In retrieval, the goal is to return the $k$ closest points in $X$ to a query $q\in\mathbb{R}^{d}$, i.e., $\mathcal{N}_{k}(q)=\arg\min_{\begin{subarray}{c}S\subset X,|S|=k\end{subarray}}\sum_{x\in S}\|q-x\|^{2}$, typically under Euclidean or cosine distance [2]. For classification, each point $x\in X$ carries a label $y_{x}\in\mathcal{Y}$, and the query’s label is predicted via majority vote among neighbors: $\hat{y}(q)=\arg\max_{c\in\mathcal{Y}}\sum_{x\in\mathcal{N}_{k}(q)}\mathbb{I}[y_{x}=c]$, where $\mathbb{I}[\cdot]$ is the indicator function [12]. Alternatively, a learned decision function $f:\mathbb{R}^{d}\to\mathcal{Y}$ can be trained directly over the embedding space [23]. In both cases, the density, structure, and discriminative power of $X$ are crucial to performance.
Approaches that build a static index, such as hashing techniques [25], and vector quantization methods [19], which partition a set of data points $X$, construct the index independently of the data distribution. Rather than fitting the indexing structure to the spatial and geometric properties of set of data point $X$, these methods project the data points to a pre-defined and fixed structure. Such methods usually impose generic assumptions on the distribution of $X$, to provide theoretical error bound guarantees. Partitions based on such assumptions often misalign with underlying high-dimensional data distributions, which is also known as the curse of clustering, while distance metrics fall short due to curse of dimensionality.
Several ANN methods build indices aligned with the data distribution, including LSH variants [14, 5], quantization-based approaches [19, 15], clustering methods [6], and tree-based partitioning [31]. These techniques aim to accelerate retrieval by organizing the space into regions aligned with approximate neighborhood structure. While effective in certain regimes, they exhibit key limitations: high training cost for large-scale data, fragile performance under distribution shift, and rigid, non-adaptive structure post-training. In contrast, deep learning methods model the data manifold directly, capturing local density and semantic continuity to enable flexible, unified retrieval and classification [9, 45].
This begs the question: How can we efficiently learn the high dimensional space to perform fast, accurate and semantic approximate nearest neighbor searches? In other words can we learn a geometric representation $\mathcal{G}_{\theta}$ of $\mathbb{R}^{d}$, parameterized by, e.g., Gaussians or neural fields, that enables efficient, robust solutions to both retrieval and classification, while adapting to the structure and distribution of $X$?
@@ 1.1 Related Work
Traditional ANNS approaches. Approximate nearest neighbor search (ANNS) methods traditionally fall into three categories: (i) Indexing, including LSH-based hashing [25, 19, 46], randomized projections and Johnson–Lindenstrauss-based embeddings [4], vector quantization [20, 26, 3], and tree-based methods [28]; (ii) Sketching, which compresses data via hash functions to accelerate approximate distance computations [16, 35, 36, 34, 13, 27, 43]; and (iii) Graph-based search, which performs greedy traversal on structures such as Navigating Spreading-out Graphs [18] or Hierarchical Navigable Small World Graphs [38]. While effective in practice, these methods typically rely on fixed partitions or hand-crafted heuristics that do not adapt to the geometry of learned embedding spaces.
The rise of high-dimensional embeddings from deep and geometric representation learning has shifted the focus of ANNS toward learning-based approaches. In contrast to classical features, these embeddings are structured: they reflect semantic similarity, lie on low-dimensional manifolds, and exhibit highly nonuniform density. This calls for indexing methods that do not impose rigid, predefined partitions, but instead learn representations that align with the intrinsic geometry of the data.
Learning-based indexing. A prominent direction is learning to index, where models are trained to build adaptive partitions over the data [14, 21]. Neural LSH [14], for example, learns a partitioning of $\mathbb{R}^{d}$ via a classifier trained on a balanced $k$-NN graph. However, such balance constraints often conflict with real data distributions—especially under long-tailed or clustered regimes—leading to poor locality. Beyond hard partitioning, recent approaches have explored classification-based routing [24] and neural scoring of clusters [11], aiming to improve query-time ranking. Yet these methods still rely on discretized, often rigid space decompositions. In contrast, we advocate for a geometric learning perspective: learning smooth, structure-aware representations that are density and semantic aligned. Such representations enable unified treatment of retrieval and classification, generalize under distribution shift.
Contributions. We propose GARLIC, a method for learning geometry-aware representations of high-dimensional spaces via a set of Gaussians optimized for both approximate nearest neighbor retrieval and classification. GARLIC learns a probabilistic representation that adapts to local data density and semantic similarity, aligning the shape and placement of Gaussians with the underlying structure of the data. In particular, our method adapts to regions with varying intrinsic dimensionality (e.g., curves, surfaces, volumes), enabling each Gaussian to act as an informative geometric primitive and local manifold approximator, rather than a static partitioning cell.
  • We introduce a geometric representation learning approach inspired by Gaussian Splatting [30], that implicitly partitions the input space, optimized for both retrieval and classification using a unified training objective.
  • We develop adaptive refinement strategies (split, clone) tailored for high-dimensional spaces, and a localized quantization method for handling density variation, improving search efficiency and robustness w.r.t. noise.
  • We design a novel information-theoretic objective combining Mahalanobis-based divergence, entropy-reducing confidence, and geometry-preserving anchoring, aligning the Gaussian structure with the data manifold.
  • We benchmark GARLIC for retrieval and classification. Our approach exhibits state-of-the-art performance on both tasks w.r.t. runtime-recall tradeoff, especially for small candidate sets. The source code of our approach will be publicly available.
@@ 2 Method
GARLIC uses a set of $d$-dimensional Gaussians $\mathcal{G}=\{\mathcal{N}(\mu_{i},\Sigma_{i})\}_{i}$, that adapt to the underlying data distribution, provided as a set of points $x\in X\subset\mathbb{R}^{d}$, to represent high-dimensional space for efficient ANNS and classification. We choose Gaussians because they capture both local geometry and statistical structure, unlike simpler volumes such as boxes or zonotopes, and express anisotropy, curvature, and density variation in a unified, continuous form. The closed form Mahalanobis scoring and differentiability also make them well-suited for end-to-end optimization. Starting with data points, we represent the space with the set of Gaussians $\mathcal{G}$ parameterized by means and covariance matrices via Cholesky decomposition (Sec. 2.1). We optimize the coefficients (mean and cholesky) of the Gaussians using objectives that aim to balance coverage, point assignment confidence, and structural consistency (Sec. 2.2). Unlike traditional partitioning methods that enforce or tend towards balanced divisions [14, 19, 6, 31, 1, 41, 39], our approach adaptively refines the set of Gaussians $\mathcal{G}$ through splitting and cloning operations, creating a data-dependent partitioning that accommodates varying densities (Sec. 2.3). To further enhance search efficiency, i.e., narrow down the number of searches, a local hyper-spherical quantization index within each Gaussian bucket (Sec. 2.4) is built. At query time, our method uses Mahalanobis distances for fast bucket selection and prioritized bin search for efficient retrieval (Sec. 2.5). This combination of techniques allows GARLIC to effectively capture the natural clustering structure of high-dimensional data while maintaining search performance, as illustrated in Fig. 1.
Figure 1: Overview of GARLIC. The method starts with input vector points that are used to initialize a set of Gaussians parameterized by mean $\boldsymbol{\mu}$ and covariance matrices via Cholesky decomposition. During optimization, these Gaussians are refined through our novel objectives that balance coverage, assignment confidence, and structural consistency. The representation is adaptively improved through split and clone operations creating a data-dependent semantic partitioning that better reflects varying data densities in different regions of the feature space.
@@ 2.1 Gaussian Parameterization
Each Gaussian $g_{i}$ is defined by its mean $\boldsymbol{\mu_{i}}\in\mathbb{R}^{d}$ and covariance matrix $\boldsymbol{\Sigma}_{i}\in\mathbb{R}^{d\times d}$, parameterized as $\boldsymbol{\Sigma}_{i}=\mathbf{L}_{i}{\mathbf{L}_{i}}^{\top}$, $\mathbf{L}_{i}\in\mathbb{R}^{d\times d}$ being a lower triangular Cholesky factor of matrix $\boldsymbol{\Sigma}_{i}$, ensuring positive semi-definiteness (a requirement for valid Gaussian distributions). The means $\boldsymbol{\mu}_{i}$ are initialized using Κ-Μeans++ [7] on a subset of the training set. Cholesky factors $\mathbf{L}_{i}$ are initialized as:
$\mathbf{L}_{i}=\log(\bar{\delta_{\mathcal{E}}})\cdot\mathbf{I}_{d}+\boldsymbol{\epsilon},$
where $\mathbf{I}_{d}\in\mathbb{R}^{d\times d}$ is the identity matrix, and $\bar{\delta_{\mathcal{E}}}$ is the mean Euclidean distance from each $\boldsymbol{\mu}_{i}$ to its three nearest neighbors. This ensures that the initial scale of each Gaussian reflects the local data density. The perturbation $\boldsymbol{\epsilon}\in\mathbb{R}^{d\times d}$ is a random lower-triangular matrix, defined below:
$\boldsymbol{\epsilon}_{jk}=\begin{cases}2\cdot\sigma(r_{jk})-1,&j>k\\ 0,&\text{otherwise}\end{cases}$
where $r_{jk}\sim\mathcal{U}(0,0.01)$ and $\sigma(\cdot)$ is the sigmoid function. This results in diagonal-dominant $\mathbf{L}_{i}$, which stabilizes the optimization process while enabling flexible, anisotropic covariance structures. This initialization strategy attempts to align the initial Gaussian scales with the local data density and to ensure sufficient diversity among Gaussians for adaptive refinement in later stages.

### v2

@@ Abstract
We present GARLIC, a representation learning approach for Euclidean approximate nearest neighbor (ANN) search in high dimensions. Existing partitions tend to rely on isotropic cells, fixed global resolution, or balanced constraints, which fragment dense regions and merge unrelated points in sparse ones, thereby increasing the candidate count when probing only a few cells. Our method instead partitions $\mathbb{R}^{d}$ into anisotropic Gaussian cells whose shapes align with local geometry and sizes adapt to data density. Information-theoretic objectives balance coverage, overlap, and geometric alignment, while split/clone refinement introduces Gaussians only where needed. At query time, Mahalanobis distance selects relevant cells and localized quantization prunes candidates. This yields partitions that reduce cross-cell neighbor splits and candidate counts under small probe budgets, while remaining robust even when trained on only a small fraction of the dataset. Overall, GARLIC introduces a geometry-aware space-partitioning paradigm that combines information-theoretic objectives with adaptive density refinement, offering competitive recall–efficiency trade-offs for Euclidean ANN search.
@@ 1 Introduction
Let $X=\{x_{i}\}_{i=1}^{n}\subset\mathbb{R}^{d}$ be a finite point set, $q\in\mathbb{R}^{d}$ a query, and $\delta_{\mathcal{E}}:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$ the Euclidean distance. For an integer $k\geq 1$, the exact $k$-Nearest Neighbor Search (NNS) problem returns the $k$ closest points $N_{k}(q)\subseteq X$. Its approximate variant, $k$-ANN, relaxes this by requiring $A_{k}(q)\subseteq X$, $|A_{k}(q)|=k$, such that $\max_{a\in A_{k}(q)}\delta_{\mathcal{E}}(q,a)\leq c\,\delta_{\mathcal{E}}\!\big(q,x_{(k)}(q)\big),$ for some approximation factor $c\geq 1$, where $x_{(k)}(q)$ denotes the $k$-th true neighbor of $q$. We restrict attention to Euclidean spaces of hundreds of dimensions, and to indices defined by partitions of $\mathbb{R}^{d}$ into cells, where a query inspects only a few cells and re-ranks the resulting candidates. Nearest neighbor search in this setting is a canonical problem of high-dimensional geometry and algorithms, with consequences across information retrieval, computer vision, robotics, and data analysis (Lowe, 2004, Cai et al., 2021, Shakhnarovich et al., 2008, Aumüller et al., 2020, Douze et al., 2024).
ANN algorithms attempt to reduce cost in two independent ways. Sketch-based techniques (Razenshteyn, 2017, Wang et al., 2014) attempt to compress every point into a short-coded representation, a summary, so that approximate distances can be quickly evaluated. Index-based methods (Gani et al., 2016) pre-partition the space, and examine only a subset of the point set, at query time. The two approaches are complementary and often combined in practice. The focus of this work is on indexing methods, and most specifically space-partitioning in the $\bigl(\mathbb{R}^{d},\delta_{\mathcal{E}}\bigr)$ metric space (Aumüller et al., 2020). The space is divided into cells $\mathcal{B}_{g}$, each storing data points, and a query touches only those stored in the cell that contains the query point (plus a few neighbors for higher recall).
Space-partition indices are practical and efficient with their small space overhead, each cell stores a representative, e.g., a centroid, and a list of point IDs, far less than e.g., graph indices need (Malkov & Yashunin, 2018). Cells can be queried in parallel by different cores, read in one-shot by GPUs, and fetched as one block (I/O call) from the disk storage (Johnson et al., 2019, Douze et al., 2024, Jayaram Subramanya et al., 2019). These strengths hold only when the cells are well built, whether by fixed, data-independent rules or by partitions learned from the data. What drives performance is how cells are built. Data-independent (Andoni & Indyk, 2008, Andoni et al., 2018) rules fix splits a priori (e.g., random hyperplanes, simple trees), so they build in $\mathcal{O}(|X|d)$ time but ignore the geometry of $X$, and recall degrades on clustered or curved regions. Data-dependent schemes (e.g., k-means/IVF families (Jegou et al., 2010), learned hashing (Wang et al., 2015)) fit cell parameters to $X$ and usually improve recall per number of candidates visited.
In practice, partitions often tend to be isotropic, for example Voronoi cells around $k$-means centroids (Lloyd, 1982, Arthur & Vassilvitskii, 2007), and a single global number of cells $K$ is chosen. These design choices then cause predictable errors on heavy-tailed data (Clauset et al., 2009). Partition resolution, a single global number of cells $K$ applied everywhere, means dense regions get fragmented into many small (near-spherical) cells, while sparse regions are covered by a few large ones (Du et al., 1999). Balanced partitions on non-uniform data create a complementary problem: in dense zones they split true neighbors across cells, and in sparse zones they pack unrelated points together to meet the size target (Malinen & Fränti, 2014, Aumüller et al., 2020). Neighborhoods are approximately Euclidean only locally; large cells merge unrelated regions, while overly small ones fragment continuous neighborhoods. To recover locality one must either increase $K$ or probe (touch), many adjacent cells (Johnson et al., 2019, Lv et al., 2007), which raises candidate (distance) counts and hurts low-probe regimes. This leaves a concrete gap: partitions that capture local geometry and adapt their local resolution (effective cell size / expected cell cardinality) to density under a principled objective that balances reconstruction fidelity against candidate count.
Contributions. We propose GARLIC, a geometry-aware space-partition index for Euclidean ANN, optimized under an information-theoretic objective that balances coverage, overlap, and budget efficiency. Under this objective, GARLIC learns a probabilistic partition of $\mathbb{R}^{d}$ into Gaussian cells whose shape and placement align with local principal directions and whose sizes adapt to local density, adding capacity only where needed through local adaptive refinement. The resulting partition reduces cross-cell neighbor splits and candidate counts under small probe budgets.
  • Anisotropic, density-adaptive partition. GARLIC represents $\mathbb{R}^{d}$ with Gaussian cells that follow local geometry and adapt to density, improving within-cell neighbor cohesion under small candidate budgets. (Section 2.1 – 2.3)
  • Information-theoretic objective. We balance coverage, overlap, and probe efficiency via expected Mahalanobis coverage, an assignment-entropy penalty, and geometric anchoring regularization. (Section 2.2)
  • Local adaptive refinement. We add Gaussians only where needed through split/clone operations triggered by cell size and spill ratio, avoiding a single global resolution (one $K$ everywhere). (Section 2.3)
  • Budget-centric evaluation and analysis. We report competitive performance across multiple accuracy and distortion metrics under candidate and distance budgets on standard Euclidean benchmarks, and provide build/query/space complexity together with ablations that isolate each component’s contribution (Section 3, Appendix A.1).
@@ 1.1 Related Work
Traditional ANN Families.
ANN methods fall into three main families: 1. sketching and compression, which encode vectors into compact codes for fast distance estimation (e.g., product quantization (Jegou et al., 2010), optimized PQ (Ge et al., 2013), iterative quantization (Gong et al., 2012)) 2. index-based methods, which pre-organize the dataset to reduce the number of points touched at query time (e.g., IVF (Jegou et al., 2010), PCA-trees (Sproull, 1991), randomized projections and Johnson–Lindenstrauss-based embeddings (Anagnostopoulos et al., 2018) 3. graph-based methods, which traverse neighborhood graphs during search (e.g., HNSW (Malkov & Yashunin, 2018), DiskANN (Jayaram Subramanya et al., 2019)). Within indices, our work focuses on the sub-family of space-partition indices, which balance memory efficiency with parallelizability and provide a probe-based complexity model compatible with GPU and IO acceleration.
Data-Independent Partitions.
Data-independent indices split space according to fixed random rules, ignoring the geometry of the dataset. A canonical example is hyperplane LSH, which assigns points based on the sign of random projections and can be queried more flexibly via multi-probing (Indyk & Motwani, 1998, Lv et al., 2007). These methods offer theoretical guarantees and fast build times, but their isotropic and geometry-agnostic partitions lead to poor recall on clustered or manifold-structured data. GARLIC instead learns anisotropic, density-adaptive cells aligned with the underlying data geometry.
Data-Dependent Partitions.
Classical learned indices often rely on isotropic partitions with a fixed global number of cells. $k$-Means and its inverted-file variants (IVF) assign points to centroid Voronoi cells (Lloyd, 1982, Jegou et al., 2010), while PCA-trees split recursively along principal components (Sproull, 1991). Scalable extensions include mini-batch k-means (Sculley, 2010), BIRCH, which builds a hierarchical clustering tree with compact representations (Zhang et al., 1996), and BLISS, which incrementally refines partitions for large datasets (Gupta et al., 2022). These methods are efficient, but their isotropic cells and global resolution fragment dense regions and mix unrelated points in sparse ones. Neural LSH takes a different approach by building balanced cuts of the $k$-NN graph and training a classifier to extend them to $\mathbb{R}^{d}$ (Dong et al., 2020). While this can outperform k-means in some settings, the emphasis on balance rather than geometry often splits true neighbors in dense areas and merges unrelated points in sparse areas, raising candidate counts. Gaussian mixture models (GMMs) capture local covariance through Mahalanobis metrics and soft assignments (Dempster et al., 1977, Banerjee et al., 2005). These models demonstrate the benefits of anisotropy, but they maximize likelihood rather than probe efficiency and do not refine capacity locally. GARLIC combines the strengths of these directions by learning anisotropic, density-adaptive partitions with local split/clone refinement under an information-theoretic probe-budget objective, explicitly tailored to ANN retrieval.
The remainder of this work is organized as follows: Section 2 introduces the GARLIC framework, including Gaussian parameterization, the information-theoretic optimization objective, and adaptive refinement strategies. Section 3 presents our experimental evaluation on standard Euclidean benchmark datasets, a set of crucial ablation studies, and GARLIC’s limitations. Finally, conclusions are drawn in Section 4.
@@ 2 Method
GARLIC uses a collection of Gaussians $\mathcal{G}=\{\mathcal{N}(\mu_{i},\Sigma_{i})\}_{i}$, whose means and covariances adapt to the underlying data distribution, to partition $X\subset\mathbb{R}^{d}$ into cells for ANN. We choose Gaussians because their mean $\mu$ and covariance $\Sigma$ jointly encode geometry and statistics: eigenvectors of $\Sigma$ capture local principal directions, eigenvalues control anisotropy and effective dimensionality, and the Mahalanobis distance $\delta_{M}^{2}(\mathbf{x},g_{i})=(\mathbf{x}-\boldsymbol{\mu}_{i})^{\top}\boldsymbol{\Sigma}_{i}^{-1}(\mathbf{x}-\boldsymbol{\mu}_{i})=\|\mathbf{L}_{i}^{-1}(\mathbf{x}-\boldsymbol{\mu}_{i})\|_{2}^{2}$, is the canonical quadratic form associated with the covariance. Level sets of $\delta_{M}$ correspond to $\chi^{2}_{d}$ probability contours, giving a calibrated notion of coverage. Parameterizing $\boldsymbol{\Sigma}$ via its Cholesky factor yields closed-form (Section 2.1), differentiable gradients while guaranteeing positive definitenes, which makes Gaussians uniquely well-suited among partitioning primitives for end-to-end optimization, unlike boxes or zonotopes that lack smooth probabilistic distance functions. Compared to alternatives such as Gaussian log-likelihood, which adds normalization terms unrelated to retrieval efficiency, or KL divergence, which compares distributions rather than points, Mahalanobis provides a direct, efficient, and anisotropy-aware metric for both training and retrieval.
We train the Gaussians with information-theoretic objectives that balance coverage, assignment confidence, and structural consistency (Section 2.2), and refine capacity only where needed via split and clone operations, avoiding uniform or balanced partitions that fragment dense regions or merge sparse ones (Dong et al., 2020, Gong et al., 2012, Arthur & Vassilvitskii, 2006, Kumar et al., 2008, Abdullah et al., 2014, Sproull, 1991, McNames, 2001) (Section 2.3). At query time, each Gaussian cell is equipped with a local hyper-spherical quantization index to narrow candidate searches (Section 2.4), and Mahalanobis distances are again used for cell selection and prioritized bin search (Section 2.5). This approach enables GARLIC to adapt to data geometry while maintaining efficient retrieval, as illustrated in Figure 1.
[…]
Figure 1: Overview of GARLIC. Input vectors are represented by Gaussian cells (means, covariances) optimized with information-theoretic objectives for coverage, confidence, and consistency. Cell density is refined adaptively via split/clone operations, and queries use spherical quantization with Mahalanobis distance for cell selection and retrieval. Blue arrows indicate transferring information of gradients (back-propagation) or statistics.
@@ 2.1 Gaussian Parameterization
Each Gaussian $g_{i}$ is defined by its mean $\boldsymbol{\mu_{i}}\in\mathbb{R}^{d}$ and covariance $\boldsymbol{\Sigma}_{i}\in\mathbb{R}^{d\times d}$, parameterized as $\boldsymbol{\Sigma}_{i}=\mathbf{L}_{i}{\mathbf{L}_{i}}^{\top}$, $\mathbf{L}_{i}\in\mathbb{R}^{d\times d}$ being a lower triangular Cholesky factor of matrix $\boldsymbol{\Sigma}_{i}$, ensuring positive definiteness (a prerequisite for valid Gaussian distributions). The means $\boldsymbol{\mu}_{i}$ are initialized using K-Means++ (Arthur & Vassilvitskii, 2007) on a subset of the training set. Cholesky factors $\mathbf{L}_{i}$ are initialized as: $\mathbf{L}_{i}=\log(\bar{\delta_{\mathcal{E}}})\cdot\mathbf{I}_{d}+\boldsymbol{\epsilon},$ where $\mathbf{I}_{d}\in\mathbb{R}^{d\times d}$ is the identity matrix, and $\bar{\delta_{\mathcal{E}}}$ is the mean Euclidean distance of each $\boldsymbol{\mu}_{i}$ to its three nearest neighbors. This ensures that the initial scale of each Gaussian reflects the local data density. Perturbation $\boldsymbol{\epsilon}\in\mathbb{R}^{d\times d}$ is a random lower-triangular matrix with entries $\epsilon_{jk}=2\sigma(r_{jk})-1$ if $j>k$ and 0 otherwise, where $r_{jk}\sim\mathcal{U}(0,0.01)$ and $\sigma(\cdot)$ is the sigmoid function. This construction yields diagonal-dominant $\mathbf{L}_{i}$, stabilizing optimization while allowing anisotropic covariances.